a car moving with a velocity 10m/s is followed by another car moving with velocity 5 m/s accelerating with 5m/s^2 and is 15m from the first vehicle.when will the two vehicles cross each other?(please solve it for class 9 level with steps)
1
Expert's answer
2018-06-17T08:22:08-0400
Let t is the time when cars will meet. The displacement of the cars
l_1=v_1 t=10t
l_2=v_2 t+(at^2)/2=5t+2.5t^2
Since
l_2-l_1=15 m
We get
5t+2.5t^2-10t=15
t^2-2t-6=0
The positive root
t=1+√7=3.65 s